OpenStudy (anonymous):

sequence is : 1, 2, 3, 5, ?, 13, 21, ?

OpenStudy (anonymous):

fibbonacci sequence

OpenStudy (anonymous):

find a digit by adding the previoustwo digits

OpenStudy (anonymous):

8 and 44

OpenStudy (anonymous):

8 and 34

OpenStudy (anonymous):

8 and 34 are the answer :)

OpenStudy (anonymous):

it is like 1+2 = 3 3+2 = 5 5+3 = 8 8+5 = 13 13+8 = 21 21+13 = 44

OpenStudy (anonymous):

cool! needed that for a job interview! weird!

OpenStudy (anonymous):

21 + 13 is 34 ;)

OpenStudy (anonymous):

i m sorry 34

OpenStudy (anonymous):

8 and 34 it is , thanks ;)

OpenStudy (anonymous):

you are welcome :) Good luck on the job interview